# Optimizing the Training Schedule of Multilingual NMT using Reinforcement Learning

### v1

@@ 1 Introduction
Multilingual neural machine translation systems (NMT) are particularly effective to enable the translation of low-resource languages (LRLs) when these are accompanied, in the training data, by related high-resource languages (HRLs) (Gu et al., 2018; Neubig and Hu, 2018). Including HRLs in the training data reduces the chance of overfitting to the LRLs and improves the quality of their translation.
@@
We propose to use reinforcement learning (RL) to optimize the training schedule of many-to-one NMT systems, i.e. to improve the training process and the resulting system compared to a fixed sampling strategy. We enable our systems to select the source language of the batch at each training step, based on a learned estimate of the model’s competence on each language in terms of its loss on a development set. Unlike fixed strategies, such as training on the hardest language, we leverage RL to let the model find better adaptive strategies.
@@
  • Based on the Deep Q Network algorithm (DQN) (Mnih et al., 2013), we design a RL-based model in which the expected rewards are generated by an auxiliary neural network that is trained in parallel with the NMT system.
@@
  • DQN outperforms (in terms of BLEU and COMET scores) previous training schedules used for multilingual NMT: monolingual minibatches sampled in proportion of each language or equally, or multilingual ones.
@@ 2 Related Work
When the teacher model is in charge of the training schedule of the student, it may use reinforcement learning (RL), with the student model playing the role of the environment (Shen and Zhao, 2024). RL has proved particularly useful at training large language models to follow instructions (Ouyang et al., 2022), initially using PPO (Schulman et al., 2017) and then other algorithms (Rafailov et al., 2023; Ethayarajh et al., 2024), but these methods are not designed to optimize training schedules. While it is possible to use curriculum learning to train RL-based models (Narvekar et al., 2020), e.g. by presenting them with increasingly difficult problems, we focus here on the use of RL to train a teacher model, in the field of multilingual NMT.
@@
Neubig and Hu (2018) study the upsampling of the LRL data when building minibatches, and observe that keeping the original proportions of HRL and LRL performs marginally better. However, Johnson et al. (2017) and Aharoni et al. (2019) sample each batch uniformly from a concatenation of all language pairs. Arivazhagan et al. (2019) compare a simple concatenation with uniform balancing, and observe better results for LRLs when using temperature-based upsampling, which has been favored afterwards (Conneau et al., 2020; Tang et al., 2021). The translation capabilities of large language models (LLMs) have also been explored: Zhu et al. (2024) compared several recent LLMs, and showed that they can achieve state-of-the-art results when translating HRLs, but highlighted their limitations in translating LRLs compared to NMT models.
@@
Jean et al. (2019) compare adaptively upsampling a language depending on various criteria, observing best results on LRLs when dynamically changing the norm of the gradient. Wang et al. (2020b) adaptively balance the languages by learning their weights from the model’s competence on a development set. Zhang et al. (2021) design a dynamic sampling strategy which measures per-language competence but also evaluates LRL competence through a related HRL’s competence. Wu et al. (2021) also balance the data dynamically, but measure a model’s uncertainty as the variance over several runs of Monte Carlo dropout. Estimates of competence using the evolution of the loss of the NMT system have been proposed by Zaremoodi and Haffari (2019), who use its absolute value, by Xu et al. (2020), who use its relative decrease, and by Atrio et al. (2024), who use KL divergence between consecutive states of the weights of an entire Transformer network.
@@ 3 Two RL Algorithms for Optimizing Training Schedules
In this study, as we use for training only monolingual batches, the set of possible actions $\mathscr{A}$ is simply the set of source languages. The states $\mathscr{S}$ of the system are the values of the parameters of the neural network and of the optimizer. However, these are too numerous to be sensibly observed at each step. Drawing inspiration from Wu et al. (2018a), we compute the current state of the model as the vector of cross-entropy loss values obtained from the NMT system over a development batch of sentences. [^2] Alternatively, MT-specific metrics such as BLEU or COMET could be used instead of the cross-entropy loss, but computing them is more costly and therefore we do not use them in this study. We use the score $X_{t}$ of the model at time step $t$ to compute the reward $R_{t}$ as the decrease of the loss of the NMT system between the last two time steps: $R_{t}=X_{t}-X_{t-1}$. The loss values are computed on a development minibatch of data selected from the current language, or, in some experiments, on a multilingual minibatch.
@@ 5.2 Translation Performance
Figure 3: Proportions of data seen during training, as optimized by the TSCL and DQN algorithms, in comparison to uniform (1/8) or to original proportions.
@@ 5.4 Role of Hyperparameters
For TSCL, we observe first that adding LRLs during warm-up (with uniform frequencies), or skipping warm-up entirely (thus starting with the highest learning rate), are not good options (2nd and 3rd lines of Table 5). Instead, cross-lingual transfer from HRLs to LRLs becomes fully beneficial only with a 16k step warm-up on HRLs. Moreover, convergence is twice slower without warm-up. The smoothing coefficient $\alpha$ can vary around the default value of 0.3 with a small decrease in performance ($\alpha=0.1$ is shown in the 4th line) and so can $\epsilon$ for the $\epsilon$-greedy policy ($\epsilon=0.3$ instead of 0.1 is shown in the 5th line). Finally, whether an action is selected every 10 steps or at every step results in comparable scores.
@@ 6 Conclusion and Perspectives
In this paper, we have presented two algorithms for optimizing the training schedule of multilingual NMT models when a mixture of HRLs and LRLs must be learned on the source side. The TSCL algorithm models the expected return of each action by smoothing past observations, while DQN trains a neural network to perform this estimation and to select the optimal action.
@@
The relevance of our proposal should be tested with additional datasets combining HRLs and related LRLs, and with other neural architectures for which cross-lingual transfer is key to ensure acceptable performance on LRLs. In such cases, an optimized training schedule across available resources is also likely to be beneficial.
@@ Acknowledgments
We are grateful for its support to the Swiss National Science Foundation (DOMAT grant n. 175693, On-demand Knowledge for Document-level Machine Translation and EXOMAT grant n. 228494, External Knowledge for Low-resource Machine Translation). We thank Giorgos Vernikos for feedback on earlier versions of the paper.

### v2

@@ Abstract
Optimizing the Training Schedule of Multilingual NMT
using Reinforcement Learning
Alexis Allemann1, Àlex R. Atrio2,*, Andrei Popescu-Belis1,3 1HEIG-VD / HES-SO, 1401 Yverdon-les-Bains, Switzerland 2Pi School, Rome, Italy 3EPFL, 1015 Lausanne, Switzerland Correspondence: andrei.popescu-belis@heig-vd.ch
footnotetext: Work performed while at HEIG-VD / HES-SO.
©  2025 The authors. This article is licensed under a Creative Commons 4.0 licence, no derivative works, attribution, CC-BY-ND.
@@ 1 Introduction
Multilingual neural machine translation (NMT)is particularly effective to enable the translation of low-resource languages (LRLs) when they are accompanied, in the training data, by related high-resource languages (HRLs) (Gu et al., 2018; Neubig and Hu, 2018). Including HRLs in the training data reduces the chance of overfitting to the LRLs and improves translation quality.
@@
We propose to use reinforcement learning (RL) to optimize the training schedule of many-to-one NMT systems, i.e. to improve the training process and the resulting system compared to a fixed sampling strategy. We enable our systems to select the source language of the batch at each training step, based on a learned estimate of the model’s competence on each language in terms of loss on a development set. Unlike fixed strategies, such as training on the hardest language, we leverage RL to let the model find better strategies.
@@
  • Based on the Deep Q Network algorithm (DQN) (Mnih et al., 2013), we design a RL-based model in which the expected rewards are generated by an auxiliary neural network trained in parallel with the NMT system.
@@
  • DQN outperforms in terms of BLEU and COMET scores previous training schedules used for multilingual NMT: monolingual minibatches sampled equally, or in proportion of each language, or multilingual batches.
@@ 2 Related Work
When the teacher model is in charge of the training schedule of the student, it may use reinforcement learning (RL), with the student model playing the role of the environment (Shen and Zhao, 2024). RL has proved particularly useful at training large language models to follow instructions (Ouyang et al., 2022), initially using PPO (Proximal Policy Optimization, Schulman et al., 2017) and then other algorithms (Rafailov et al., 2023; Ethayarajh et al., 2024), but these methods are not designed to optimize training schedules. While it is possible to use curriculum learning to train RL-based models (Narvekar et al., 2020), e.g. by presenting them with increasingly difficult problems, we focus here on the use of RL to train a teacher model, in the field of multilingual NMT.
@@
Neubig and Hu (2018) study the upsampling of the LRL data when building minibatches, and observe that keeping the original proportions of HRL and LRL performs marginally better. However, Johnson et al. (2017) and Aharoni et al. (2019) sample each batch uniformly from a concatenation of all language pairs. Arivazhagan et al. (2019) compare simple concatenation with uniform balancing, and observe better results for LRLs when using temperature-based upsampling, which was favored afterwards (Conneau et al., 2020; Tang et al., 2021).
The translation capabilities of large language models (LLMs) have also been explored: Zhu et al. (2024) compares several recent LLMs and shows that they can achieve state-of-the-art results when translating HRLs, but highlights their limitations in translating LRLs compared to NMT models. One of the leading open-weights LLMs for MT, Tower Instruct (Alves et al., 2024), is fine-tuned on a large set of translation-related tasks in 10 HRLs, with no particular scheduling of the fine-tuning data, and no reinforcement learning.
@@
Jean et al. (2019) compare adaptively upsampling a language depending on various criteria, observing best results on LRLs when dynamically changing the norm of the gradient. Wang et al. (2020b) adaptively balance the languages by learning their weights from the model’s competence on a development set. Zhang et al. (2021) design a dynamic sampling strategy which measures per-language competence but also evaluates LRL competence through a related HRL’s competence. Wu et al. (2021) also balance the data dynamically, but measure a model’s uncertainty as the variance over several runs of Monte Carlo dropout. Estimates of competence using the evolution of the loss of the NMT system have been proposed by Zaremoodi and Haffari (2019), who use its absolute value, by Xu et al. (2020), who use its relative decrease, and by Atrio et al. (2024), who use Kullback-Leibler divergence between consecutive states of the weights of an entire Transformer network.
@@ 3 Two RL Algorithms for Optimizing Training Schedules
In this study, as we use for training only monolingual batches, the set of possible actions $\mathscr{A}$ is simply the set of source languages. The states $\mathscr{S}$ of the system are the values of the parameters of the neural network and of the optimizer. However, these are too numerous to be sensibly observed at each step. Drawing inspiration from Wu et al. (2018a), we compute the current state of the model as the vector of cross-entropy loss values obtained from the NMT system over a development batch of sentences. [^2] Alternatively, MT-specific metrics such as BLEU or COMET could be used instead of the cross-entropy loss, but computing them is more costly, therefore we do not use them here. We use the score $X_{t}$ of the model at time step $t$ to compute the reward $R_{t}$ as the decrease of the loss of the NMT system between the last two time steps: $R_{t}=X_{t}-X_{t-1}$. The loss values are computed on a development minibatch of data selected from the current language, or, in some experiments, on a multilingual minibatch.
@@ 5.2 Translation Performance
Figure 3: Proportions of data seen during training, as optimized by the TSCL and DQN algorithms, in comparison to uniform (1/8) or original proportions.
@@ 5.4 Role of Hyperparameters
For TSCL, we observe first that adding LRLs during warm-up (with uniform frequencies), or skipping warm-up entirely (thus starting with the highest learning rate), are not good options (second and third lines of Table 5). Instead, cross-lingual transfer from HRLs to LRLs becomes fully beneficial only with a 16k step warm-up on HRLs. Moreover, convergence is twice slower without warm-up. The smoothing coefficient $\alpha$ can vary around the default value of 0.3 with a small decrease in performance ($\alpha=0.1$ is shown in the 4th line) and so can $\epsilon$ for the $\epsilon$-greedy policy ($\epsilon=0.3$ instead of 0.1 is shown in the 5th line). Finally, whether an action is selected every 10 steps or at every step results in comparable scores.
@@ 6 Conclusion and Perspectives
In this paper, we presented two algorithms for optimizing the training schedule of multilingual NMT models when a mixture of HRLs and LRLs must be learned on the source side. The TSCL algorithm models the expected return of each action by smoothing past observations, while DQN trains a neural network to perform this estimation and to select the optimal action.
@@
The relevance of our proposal should be tested with additional datasets combining HRLs and related LRLs, and with other neural architectures for which cross-lingual transfer may be important to ensure acceptable performance on LRLs, particularly LLMs fine-tuned on translation tasks (Alves et al., 2024). In such cases, an optimized training schedule across available resources may also be beneficial.
@@ Acknowledgments
We are grateful for its support to the Swiss National Science Foundation, through the DOMAT grant n. 175693, On-demand Knowledge for Document-level Machine Translation and the EXOMAT grant n. 228494, External Knowledge for Low-resource Machine Translation. We thank Giorgos Vernikos for feedback on earlier versions of the paper.